OpenStudy (anonymous):

How do you solve this? Please explain I have a test tomorrow and I don't get how to solve this: 2/3 ( x-18)=-6

OpenStudy (mew55):

distribute the 2/3 so u will get 2/3x-12=6 add 12 to both sides 2/3x=18 then divide by 2/3 x=27

OpenStudy (anonymous):